Given below is a demand schedule for pizza in a market

Average Price of Pizza

Quantity demanded (millions per week)

$21

8

$18

14

$15

20

$12

26

$9

32

  1. What is the slope of the demand curve?
  1. What is the price elasticity of demand at 12$ and $6? What observation(s) can you make based on these values?

  1. Using the formulae that we used in class, how many units can you sell at $16.25?
Answer #1

A) As we can derive from the given dataset, a $3 change in price leads to a 6 unit change in quantity demanded.

Thus, slope of demand curve = -3/6 = -0.5

B) When P = 12, Q = 26

When P = 6, Q = 26+6+6 = 38

P elasticity of demand = (Change in demand/Change in P)(Average P / Average Q)

Price elasticity of demand = (-12/6)(9/32)

Price elasticity of demand = -0.5625

Since P elasticity of demand is less than 1, it means demand is inelastic in nature

C) Deriving demand curve from given dataset,

Q = 50-2P

Thus, when P = 16.25,

Q = 50-2(16.25)

Q = 17.5 units